# Auditing and Generating Synthetic Data with Controllable Trust Trade-offs

### v3

@@ Abstract
Real-world data often exhibits bias, imbalance, and privacy risks. Synthetic datasets have emerged to address these issues. This paradigm relies on generative AI models to generate unbiased, privacy-preserving data while maintaining fidelity to the original data. However, assessing the trustworthiness of synthetic datasets and models is a critical challenge. We introduce a holistic auditing framework that comprehensively evaluates synthetic datasets and AI models. It focuses on preventing bias and discrimination, ensures fidelity to the source data, assesses utility, robustness, and privacy preservation. We demonstrate the framework’s effectiveness by auditing various generative models across diverse use cases like education, healthcare, banking, and human resources, spanning different data modalities such as tabular, time-series, vision, and natural language. This holistic assessment is essential for compliance with regulatory safeguards. We introduce a trustworthiness index to rank synthetic datasets based on their safeguards trade-offs. Furthermore, we present a trustworthiness-driven model selection and cross-validation process during training, exemplified with "TrustFormers" across various data types. This approach allows for controllable trustworthiness trade-offs in synthetic data creation. Our auditing framework fosters collaboration among stakeholders, including data scientists, governance experts, internal reviewers, external certifiers, and regulators. This transparent reporting should become a standard practice to prevent bias, discrimination, and privacy violations, ensuring compliance with policies and providing accountability, safety, and performance guarantees.
@@ Introduction
Generative models have demonstrated impressive results in synthesizing high-quality data across multiple modalities from tabular and time-series data [1, 2, 3] to text [4, 5], images [6, 7, 8, 9] and chemistry [10]. We are entering a new era in training AI models, where synthetic data can be used to augment real data [11], or as a complete replacement, in the most extreme case [12]. One of the main motivations behind controllable synthetic data usage in training AI models is its promise to synthesize privacy-preserving data that enables safe sharing without putting the privacy of real users and individuals at risk. This has the potential to circumvente cumbersome processes that are at the heart of many highly regulated fields such as financial services [13] and healthcare, for example [14, 15, 16, 11]. Another motivation comes from controlling the generation process in order to balance the training data and reduce biases against protected groups and sensitive communities [17]. Finally, synthetic data also offers new opportunities in simulating non-existing scenarios, providing grounding for causal inference via the generation of counterfactuals that would help explain some observations in the absence of real data [18].
Synthetic data can take different forms, ranging from seedless approaches [19, 20], which rely on knowledge bases and rule-based generation but risk biased grounding and linkage attacks, to data generated through AI models trained on real data, which may lead to memorization, privacy breaches, and bias amplification [21, 22, 17]. These AI-generated datasets can also pose legal issues related to copyright and intellectual property [23]. Both types of synthetic data need thorough auditing for safety, privacy, fairness, and utility alignment.
@@
To address these challenges, we propose a framework for auditing the trustworthiness of synthetic data that is holistic, transversal across different modalities (tabular, time-series, computer vision and natural language) and assess the uncertainty in auditing a generative model. See Figure 1 for a summary of our approach.
Our main contributions are as follows: We introduce a holistic framework for auditing the trustworthiness of synthetic data, covering key trust dimensions like fidelity, utility, privacy, fairness, and robustness.We define a trustworthiness index that evaluates synthetic data and their downstream tasks. We provide methods for controlling trust trade-offs in synthetic data during training, notably through model selection via the trustworthiness index. We instrument transformer models across multiple modalities with these control mechanisms and refer to them as TrustFormers. By applying our framework, we demonstrate that downstream tasks using trustworthiness-index driven cross-validation often outperform those trained on real data while meeting privacy and fairness requirements. Finally, our framework offers transparency templates for clear communication of the risks of synthetic data via an audit report.
@@
Formally, given a real dataset $D_{r}$ and multiple synthetic datasets $D^{j}_{s},\leavevmode\nobreak\ j=1\dots N$. The real dataset is split to training, development/validation, and testing sets as follows:
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ D_{r}=\{D_{r,\text{train}},\leavevmode\nobreak\ D_{r,\text{val}},\leavevmode\nobreak\ D_{r,\text{test}}\}$ (1)
@@
Our synthetic data auditing framework, as depicted in Figure 1, comprises several key stages: Evaluate, Aggregate, Reweight, Index by trustworthiness, Rank, and Report. It serves as a means to enhance communication among governance experts, data scientists, internal reviewers, and external regulators. The primary personas involved in our framework are outlined in Extended Data Figure 1. The key stages of our audit framework and the personas involved are explained below (more detailed explanations of trust dimension and our framework are given in the Methods Section and Supplementary Information):
@@
    Evaluate: The governance expert and data science board collaborate to establish quantitative metrics for each trust dimension, as detailed in Extended Data Table 4, ensuring adherence to socio-technical safeguards. These metrics are then assessed by data scientists.
@@
    Re-weight: Stakeholders, including internal reviewers, governance experts, and the application owner, establish a trustworthiness profile through trade-off weights $\omega$ (examples are shown in Extended Data Table 2), indicating the relative importance of trust indices for fulfilling specific requirements.
@@
    Index by Trustworthiness: Trust dimension indices are re-weighted by the trustworthiness profile weights $\omega$ and combined via a geometric mean to produce the trustworthiness index ($\tau_{\text{trust}}(\omega)$), a context-specific measure based on the application’s needs and trustworthiness profiles.
@@
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ R^{\alpha}_{\text{trust}}=\log(\overline{\tau_{\text{Trust}}}(\omega))-\alpha\log(\Delta_{\tau}(\omega)).$ (2)
@@
    Report: To enhance transparency, our framework provides audit report templates (as seen in Supplementary Information Section Supplementary Information G) for communicating the audit results. These audit reports can be submitted to regulators or external third-party certifiers for validation.
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{TF}(\omega,\text{n-p})\text{ for non-private training and }\text{ TF}(\omega,\text{p-}\varepsilon)\text{ for private training.}$.
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{TF}(\omega_{1},\omega_{2},\text{n-p})\text{ for non-private training and }\text{ TF}(\omega_{1},\omega_{2},\text{p-}\varepsilon)\text{ for private training.}$.
Results and Discussion
In this section, we present the training and auditing of various data generation techniques, including our TrustFormer Models, across a range of use cases encompassing tabular, time series, and natural language data (refer to Extended Data Table 1). We also conduct audits on synthetic data generated by state-of-the-art diffusion models, particularly in the computer vision domain, focusing on the ImageNet classification task and assessing their generalization under distribution shifts. The models under audit are summarized in Extended Data Table 3, and we employ our trust dimension indices to assess each model’s compliance with specific trust dimensions. For a given trustworthiness profile weights $\omega$ (refer to Extended Data Table 2 for examples and their interpretation), we rely on our trustworthiness index to conduct TrustFormer model selection and rank the models based on their alignment with predefined safeguards (given in the trustworthiness profile $\omega$).
Tabular use cases
In this Section, we showcase our auditing framework and trustworthiness index driven model selection for TrustFormer on three tabular datasets: Bank Marketing Dataset[47], Recruitment Dataset [48], and the Law School Admission Council Dataset [49]. Please refer to Extend Table 1 and Methods section for details of the datasets, downstream tasks , safeguards and TF models training details.
Audit Results
For the Bank Marketing Dataset, we present average trust dimension indices and their "variances" for non-private and private synthetic data generated by TrustFormers models (TF) in Figure 2 (a) and (b). In this case, TF models were selected across various trustworthiness profiles (as outlined in Extended Data Table 2) using the trustworthiness index from Equation (4) for $\alpha=0$. We also include results for non-private and private baselines (refer to Extended Data Table 3), shown in panels (a) and (b). Additionally, we provide information about the utility, fairness, and robustness of downstream tasks trained on real data as a reference for both panels (a) and (b). Panel (c) displays the ranking according to trustworthiness index of synthetic data across different trustworthiness profiles based on the same $\alpha$ value. Similar plots for the Recruitment and Law School datasets can be found in Extended Data Figure 2 and in Supplementary Information I Figure 8, respectively, for $\alpha=0$. To account for data split uncertainties, we’ve included audit results for $\alpha=0.1$, depicted in Figure 3 for the Law School dataset, and in the Supplementary Information for the other datasets (see Figures 9 and 10 in Supplementary Information I).
For both Bank Marketing dataset and Recruitment datasets we see that in both cases $\alpha=0$ and $0.1$, TF synthetic data outperforms other baselines across all trustworthiness profiles $\omega$ in terms of its trustworthiness index . We see in Panel (c) Figure 2 and Extended Data Figure 2, a clear phase transition between private and non private TF models depending on wether the trustworthiness profile highlights privacy as a requirement or not. This effect is achieved through our trustworthiness index model selection that offers control over trust trade-offs. On the other hand, for the law school dataset, TF models fall short with respect to baselines when uncertainty is not considered ($\alpha=0$, Supplementary Figure 8 ), but lead to top performing synthetic data when uncertainty is considered ($\alpha=0.1$, Figure 3 ). This highlights the importance of assessing the uncertainty in auditing synthetic data.
Time-Series Use Cases
Analysis of Results
We compare the performance of the two decoding strategies considered (multinomial and top-k decoding) for synthetic data generation from TF models and study how it impact trust trade-offs. Note that we used herein fixed data splits from the literature and we don’t report therefore the uncertainty of the audit. Extended Data Table 6 gives the ranking of these synthetic datasets using the trustworthiness index for all trustworthiness profiles. We see that for some trustworthiness profiles, non-private TrustFormer with multinomial decoding stands out, while top-k decoding outperforms it for other profiles. On private models we see that TF models with top-k decoding are the best at balancing the privacy/utility trade-offs. This shows that our auditing framework highlights the effect of hyper-parameters choices such as decoding strategies and their impacts on all trust dimensions
Results and Discussion
We audit the synthetic dataset sampled from different epochs (namely after 3, 5, 7 and 9) during the fine-tuning process. The downstream task we consider in this audit is the in hospital mortality prediction task, where the protected community is the ethnicity "ASIAN" [55]. As an embedding E, we use the original pre-trained BioGPT model [56] to extract embeddings for the synthetic notes as it has the capability of representing the biomedical domain. In Table 2 we see that our trustworthiness index driven model selection allows a controllable trade-off between utility and fairness : epoch 7 has a better utility fairness trade-off than the model at the last epoch that has higher utility. Therefore it is favorable to select the model at epoch 7 at the price of a reduced utility but with an enhanced fairness, which is of paramount importance for this use-case.
We consider here the use of synthetic image data for training classifiers on one of the key computer vision datasets, Imagenet [58], with a focus on the utility and generalization propreties of these classifiers on the Imagenet test set and under distribution shifts. Imagenet consists of $1.2$ Million images of size 256x256x3 and 1000 categories. Authors in [59] constructed Imagenetv2 test set, that consists of a distribution shift from the original imagenet distribution and reported a performance drop between $11\%-14\%$. Recent works [60] and [61] showed promising results of synthetic data from diffusion models (a family of generative models that uses diffusion techniques [8] ) in improving Imagenet classification. Following these promising works, we synthesize $1.2$ M labeled images from the Imagenet 256x256 pretrained guided diffusion models from OpenAI [62].Table 3 presents ResNet50 [63] performance on ImageNet-v1 and v2 under Gaussian noise, considering four training data scenarios: (1) real Imagenet; (2) synthetic data; (3) a hybrid with real and synthetic images ( real to synthetic ratio $1/0.5$ ); and (4) an equal mix totaling 2.4 million images. Results show synthetic-only models lag, but integrating with real images enhances robustness, especially in noisier settings (sigma 0.3 and 0.5). Combining synthetic data with real images improves a model’s resilience to noise and image corruption.
Conclusion

### v4

@@ Abstract
Real-world data often exhibits bias, imbalance, and privacy risks. Synthetic datasets have emerged to address these issues by enabling a paradigm that relies on generative AI models to generate unbiased, privacy-preserving data while maintaining fidelity to the original data. However, assessing the trustworthiness of synthetic datasets and models is a critical challenge. We introduce a holistic auditing framework that comprehensively evaluates synthetic datasets and AI models. It focuses on preventing bias and discrimination, ensuring fidelity to the source data, and assessing utility, robustness, and privacy preservation. We demonstrate our framework’s effectiveness by auditing various generative models across diverse use cases like education, healthcare, banking, and human resources, spanning different data modalities such as tabular, time-series, vision, and natural language. This holistic assessment is essential for compliance with regulatory safeguards. We introduce a trustworthiness index to rank synthetic datasets based on their safeguards trade-offs. Furthermore, we present a trustworthiness-driven model selection and cross-validation process during training, exemplified with “TrustFormers” across various data types. This approach allows for controllable trustworthiness trade-offs in synthetic data creation. Our auditing framework fosters collaboration among stakeholders, including data scientists, governance experts, internal reviewers, external certifiers, and regulators. This transparent reporting should become a standard practice to prevent bias, discrimination, and privacy violations, ensuring compliance with policies and providing accountability, safety, and performance guarantees.
Index Terms:
Trustworthy AI, Synthetic Data, Auditing, Generative AI.
@@ Introduction
Generative models have demonstrated impressive results in synthesizing high-quality data across multiple modalities from tabular and time-series data [88, 84, 82] to text [20, 31], images [91, 52, 100, 19] and chemistry [13]. We are entering a new era in training AI models, where synthetic data can be used to augment real data [26], or as a complete replacement, in the most extreme case [42]. One of the main motivations behind controllable synthetic data usage in training AI models is its promise to synthesize privacy-preserving data that enables safe sharing without putting the privacy of real users and individuals at risk. This has the potential to circumvent cumbersome processes that are at the heart of many highly regulated fields such as financial services [11] and healthcare, for example [29, 17, 34, 26]. Another motivation comes from controlling the generation process in order to balance the training data and reduce biases against protected groups and sensitive communities [30]. Finally, synthetic data also offers new opportunities in simulating non-existing scenarios, providing grounding for causal inference via the generation of counterfactuals that would help explain some observations in the absence of real data [38].
Synthetic data can take different forms, ranging from seedless approaches [8, 58], which rely on knowledge bases and rule-based generation but incur risks of biased grounding and linkage attacks, to data generated through AI models trained on real data, which may lead to memorization, privacy breaches, and bias amplification [24, 47, 30]. These AI-generated datasets can also pose legal issues related to copyright and intellectual property [109]. Both types of synthetic data need thorough auditing for safety, privacy, fairness, and utility alignment.
@@
To address these challenges, we propose a framework for auditing the trustworthiness of synthetic data that is holistic, transversal across different modalities (tabular, time-series, computer vision and natural language) and assess the uncertainty in auditing a generative model (see Figure 1 for a summary of our approach).
Figure 1: Summary diagram of our proposed holistic synthetic data auditing framework. For each trust dimension (fidelity, privacy, utility, fairness, and robustness), we evaluate multiple metrics on the synthetic data and quantify their uncertainty. Metrics are aggregated within each trust dimension, which results in trust dimension indices. These indices are re-weighted with desired trust trade-offs to produce the trustworthiness index. Different synthetic datasets are then ranked using this trustworthiness index, and a summary of the audit is written to an audit report. The ranking produced by our audit enables comparison of different synthetic data produced by various generative modeling techniques, and aids the model selection process for a given generation technique, allowing its alignment with prescribed safeguards. The model selection is performed via trustworthiness index driven cross-validation, which results in controllable trust trade-offs by producing new ranks for different desired weighing trade-offs for a given application and use case.
Our main contributions are as follows: we introduce a holistic framework for auditing the trustworthiness of synthetic data, covering key trust dimensions like fidelity, utility, privacy, fairness, and robustness. We define a trustworthiness index that evaluates synthetic data and their downstream tasks. We provide methods for controlling trust trade-offs in synthetic data during training, notably through model selection via the trustworthiness index. We instrument transformer models across multiple modalities with these control mechanisms and refer to them as “TrustFormers”. By applying our framework, we demonstrate that downstream tasks using trustworthiness-index-driven cross-validation often outperform those trained on real data while meeting privacy and fairness requirements. Finally, our framework offers transparency templates for clear communication of the risks of synthetic data via an audit report.
@@
Formally, given a real dataset $D_{r}$ and multiple synthetic datasets $D^{j}_{s},~{}j=1,\ldots,N$. The real dataset is split to training, development/validation, and testing sets as follows:
$D_{r}=\{D_{r,\text{train}},~{}D_{r,\text{val}},~{}D_{r,\text{test}}\}.$ (1)
@@
Figure 2: Auditing Platform and workflows connecting different stakeholders (e.g., data scientists, data governance experts, internal reviewers, external certifiers, and regulators) from model development to audit and certification via a synthetic data auditing report.
@@
Our synthetic data auditing framework, as depicted in Figure 1, comprises several key stages: Evaluate, Aggregate, Reweight, Index by trustworthiness, Rank, and Report. It serves as a means to enhance communication among governance experts, data scientists, internal reviewers, and external regulators. The primary personas involved in our framework are outlined in Figure 2. The key stages of our audit framework and the personas involved are explained below (more detailed explanations of trust dimension and our framework are given in the Methods Section and Supplementary Information):
@@
    Evaluate: The governance expert and data science board collaborate to establish quantitative metrics for each trust dimension, as detailed in Table I, ensuring adherence to socio-technical safeguards. These metrics are then assessed by data scientists.
@@
    Re-weight: Stakeholders, including internal reviewers, governance experts, and the application owner, establish a trustworthiness profile through trade-off weights $\omega$ (examples are shown in Table II), indicating the relative importance of trust indices for fulfilling specific requirements.
@@
    Index by Trustworthiness: Trust dimension indices are re-weighted by the trustworthiness profile weights $\omega$ and combined via a geometric mean to produce the trustworthiness index, $\tau_{\text{trust}}(\omega)$, a context-specific measure based on the application’s needs and trustworthiness profiles.
@@
    $R^{\alpha}_{\text{trust}}=\log(\overline{\tau_{\text{Trust}}}(\omega))-\alpha\log(\Delta_{\tau}(\omega)).$ (2)
@@
    Report: To enhance transparency, our framework provides audit report templates (as seen in Supplementary Information Section -T) for communicating the audit results. These audit reports can be submitted to regulators or external third-party certifiers for validation.
@@
[…]
@@
[…]
Dimension Metric Polarity Debiasing Tabular Time Series NLP
Fidelity Evaluated between $D_{r,train}$ and $D_{s}$:
Maximum Mean Discrepancy (MMD) SNR [64] $-1$ N/A (D/E) (E) (E)
[…]
Fréchet Inception Distance (FID)[51, 98] $-1$ N/A (D/E) (E) (E)
Precision/ Recall [65] $+1$ N/A (D/E) (E) (E)
Chi Squared $-1$ N/A (D)
$\ell_{2}$ mutual Information difference $-1$ N/A (D)
Privacy Evaluated between $D_{r,train}$ and $D_{s}$:
Exact Replicas Count $-1$ N/A (D)
k-nearest neighbor median distance [27] $+1$ N/A (D/E) (E) (E)
k-nearest neighbor mean distance [27] $+1$ N/A (D/E) (E) (E)
Utility Classifier trained on $D_{s}$
Evaluated on $D_{r,val}$ validation and $D_{r,test}$@ test:
Accuracy/ precision/ recall/ F1 score of:
Linear Logistic Regression $+1$ (D/E) (E) (E)
Nearest Neighbor classification $+1$ (D/E) (E) (E)
MLP $+1$ (D/E) (E) (E)
MLP / Adversarial debiasing [114] $+1$ (D/E) (E) (E)
MLP/ Fair Mixup [32] $+1$ (D/E) (E) (E)
Fairness Applicable to all classifiers in utility:
Evaluated on $D_{r,val}$ validation and $D_{r,test}$@ test:
Equal Opportunity Difference (absolute value)[15] $-1$
Average Odds Difference (absolute value) [15] $-1$
Equalized Odds Difference (absolute value) [15] $-1$
Robustness Applicable to all classifiers in utility:
Evaluated on $D_{r,val}$ validation and $D_{r,test}$@ test:
Adversarial Accuracy/ precision/ recall/ F1 score $+1$
Absolute Difference of Adversarial $-1$
and non adversarial utility metrics
TABLE I: Metrics and their associated polarities that are supported by our auditing framework under each trust dimension. Debiasing indicates if a utility classifier uses a debiasing technique. D indicates that the metric is computed on the data space after quantization. E indicates that the metric is computed in an embedding space. * refers to the same field values of the evaluated utility classifier. Note that our metrics are representative of each dimension and modality but are not exhaustive; other specialized metrics can be added and integrated seamlessly within our framework.
$\omega=(\omega_{f},\omega_{P},\omega_{U},\omega_{F},\omega_{R})$ Interpretation
$\omega_{all}$ $(100,100,100,100,100)/500$ Equal Importance
$\omega_{e(PU)}$ $(50,100,100,50,50)/350$ Privacy/Utility Emphasis
$\omega_{e(PUF)}$ $(50,100,100,100,50)/400$ Privacy/Utility/Fairness Emphasis
$\omega_{U}$ $(0,0,100,0,0)/100$ Utility only
$\omega_{PU}$ $(0,100,100,0,0)/200$ Privacy/Utility only
$\omega_{UF}$ $(0,0,100,100,0)/200$ Utility/Fairness only
$\omega_{e(UF)r(R)}$ $(50,50,100,100,0)/300$ Utility/Fairness Emphasis No Robustness
$\omega_{UFR}$ $(0,0,100,100,100)/300$ Utility/Fairness/Robustness only
$\omega_{UR}$ $(0,0,100,0,100)/200$ Utility/Robustness only
$\omega_{PUR}$ $(0,100,100,0,100)/300$ Privacy/Utility/Robustness only
TABLE II: Examples of weights trade offs of trust dimensions reflecting priorities in auditing synthetic data.
Use Case Dataset Modality Downstream Task Safeguards Policy Alignment Example
Banking Bank Marketing[79] Tabular Campaign prediction sensitive community (age); Fair Lending Act
user privacy; robustness
Recruitment UK Recruitment[41] Tabular Employment prediction sensitive community (ethnicity) NYC Law 144
user privacy; robustness
Education Law School Admission Tabular Admission prediction sensitive community (ethnicity); Equal Educational Opportunity Act
Council Dataset [110] user privacy; robustness
Financial Services Credit Card [8] Tabular time-series Fraud Detection user privacy Finance Regulation
Healthcare MIMIC-III[59] Tabular time-series Mortality prediction sensitive community (ethnicity); Patient Protection and Affordable Care Act
user privacy; robustness
Healthcare MIMIC-III Notes[6] NLP Mortality prediction sensitive community (ethnicity) Patient Protection and Affordable Care Act
user privacy; robustness
Visual Recognition Imagenet[6] Vision classification distribution shift Robust Generalization
TABLE III: Synthetic data use cases, safeguards and policy alignment.